OpenStudy (anonymous):

Help finding the indefinite integral of one over sqrt of 1 + sqrt x

OpenStudy (anonymous):

figure??

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{1}{\sqrt{1+\sqrt{x}}}\] And i know i'm gonna need trig functions.

OpenStudy (anonymous):

take root x = u^2

OpenStudy (anonymous):

\[x^{1/2} = u^{2} \]

OpenStudy (anonymous):

then differentiate \[x^{1/2}\] and get back to me

OpenStudy (anonymous):



OpenStudy (anonymous):

@cambrige okay I got the derivative of x^(1/2)|

OpenStudy (anonymous):

so dx= ur ans??

OpenStudy (anonymous):

My answer was 1 over 2root x. dx

OpenStudy (anonymous):

I'm lost now.

OpenStudy (anonymous):

@Concentrationalizing excuse me, if you could would you please help me with this problem?

OpenStudy (anonymous):

Consider u = 1 + \(\sqrt{x}\) du = \(\frac{1}{2\sqrt{x}}\)dx \(2\sqrt{x}du\) = dx Now, also notice from our substitution that \(\sqrt{x} = u-1\) Therefore we have \(2(u-1)du = dx\) Substituting these values into the integral gives: \[\int\limits_{}^{}\frac{ 1 }{ \sqrt{1+\sqrt{x}} }dx = \int\limits_{}^{}\frac{ 2(u-1) }{ \sqrt{u} }du\] Now we simply split this into two integrals and solve each. \[\int\limits_{}^{}\frac{ 2u-2 }{ \sqrt{u} }du = \int\limits_{}^{}\frac{ 2u }{ \sqrt{u} }du - \int\limits_{}^{}\frac{ 2 }{ \sqrt{u} }du\] And of course the left integral reduces and we have these two integrals \[2\int\limits_{}^{}\sqrt{u}du- 2\int\limits_{}^{}\frac{ 1 }{ \sqrt{u} }du\] Now just integrate each and plug back in your substitution from there. Think you can handle the integration from here? :) @study_guy2